# Stabilization of intersection Betti numbers for moduli spaces of one-dimensional sheaves on surfaces

### v1

@@ Abstract.
In this paper, we study the intersection cohomology of the moduli space of semistable one-dimensional sheaves with fixed Euler characteristic, supported in a divisor class $\beta$ on a smooth projective surface $S$. Assuming that this moduli space is irreducible, we prove that its intersection Betti numbers in a certain range are determined by a product formula derived from Göttsche’s formula for Betti numbers of Hilbert schemes of points on $S$. As an application, for Enriques and bielliptic surfaces, we show the stabilization of intersection Betti numbers of this moduli space when $\beta$ is sufficiently positive. In the Enriques case, we also prove a refined stabilization result concerning the perverse Hodge numbers when the moduli space is smooth.
@@ 1. Introduction
Throughout, we work over the field $\mathbb{C}$ of complex numbers. Let $S$ be a connected, smooth, projective surface with a fixed ample divisor $H$.
In recent years, moduli spaces of one-dimensional sheaves on $S$ have attracted considerable attention. On the one hand, their connections with enumerative geometry and moduli spaces of Higgs bundles have led to a number of new conjectures and results; see, for instance, [3, 4, 9, 14, 15, 19, 21, 28, 29, 36]. On the other hand, these moduli spaces exhibit not only asymptotic behavior analogous to that of moduli spaces of torsion-free sheaves, but also pathologies [34, §5] in contrast to the torsion-free case, revealing rich geometry that remains to be understood.
Let $\chi$ be an integer and $\beta$ a nonzero effective divisor of arithmetic genus $p_{a}(\beta)$ on $S$. Denote by $|\beta|^{\mathrm{int}}\subset|\beta|$ the open subscheme parametrizing integral curves. Consider the (coarse) moduli space $M_{\beta,\chi}$ parametrizing S-equivalence classes of one-dimensional semistable (with respect to $H$) sheaves $\mathcal{F}$} on $S$ with fixed determinant $\operatorname{det}({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and Euler characteristic $\chi({\mathcal{F}})=\chi$. Here the stability is defined in terms of the slope
$\mu({\mathcal{F}})=\frac{\chi({\mathcal{F}})}{c_{1}({\mathcal{F}})\cdot H}.$
There is a morphism
$h:M_{\beta,\chi}\to|\beta|,$
called the Hilbert–Chow morphism, which is defined by taking the Fitting support. If $|\beta|$ contains a connected smooth curve, then the morphism $h$ is an abelian fibration.
The moduli space $M_{\beta,\chi}$ is singular in general, for which the intersection cohomology $\operatorname{IH}^{*}(M_{\beta,\chi})$ (with $\mathbb{Q}$}-coefficients) might be a more suitable topological invariant than the ordinary cohomology. In this paper, we prove the following theorem about the intersection Betti numbers of $M_{\beta,\chi}$.
@@
Fix $k\in\mathbb{Z}_{\geq 0}$ and $\chi\in\mathbb{Z}$. Suppose that $M_{\beta,\chi}$ is irreducible, and let
@@
denote the codimension of $M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}})$ in $M_{\beta,\chi}$. If the following conditions hold:
@@
    $\beta$ is $\max\{2,2k-2\}$-very ample (see Definition 2.2);
@@
    $k\leq\min\{c-1,\frac{2}{3}\dim|\beta|\}$,
then we have
$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{\beta,\chi})=b_{k}^{\infty},$
where $b_{k}^{\infty}$ is defined by the following identity of formal power series
(1.1) $\sum_{k\geq 0}b_{k}^{\infty}q^{k}=\prod_{m\geq 1}\frac{(1+q^{2m-1})^{b_{1}(S)}(1+q^{2m+1})^{b_{1}(S)}}{(1-q^{2m})^{b_{2}(S)+1}(1-q^{2m+2})}.$
By Göttsche’s formula [8], the number $b_{k}^{\infty}$ is exactly the $k$-th Betti number $b_{k}(S^{[n]})$ of the Hilbert scheme $S^{[n]}$ of $n$ points for any integer $n\geq k$. The connection between Betti numbers of moduli spaces of sheaves on $S$ and $b_{k}^{\infty}$ has been observed in many studies, for example, in [5, 17] for some moduli spaces of torsion-free sheaves, in [36] for moduli spaces of one-dimensional sheaves on $\mathbb{P}^{2}$, and in [29] for those of one-dimensional sheaves on any del Pezzo surface.
Our proof of Theorem 1.1 adapts the approach in §2 and §3 of our previous joint work with Pi and Shen [29] for a more general setting. A key new ingredient in the proof is as follows.
@@
Let $\ell\geq 2$ be an integer. If $\beta$ is $(2\ell-2)$-very ample, then there exists an open subscheme $U_{\ell}\subset|\beta|^{\mathrm{int}}$ such that
@@
    $\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{\ell},|\beta|^{\mathrm{int}})\geq\ell+1$;
@@
    The open subscheme $h^{-1}(U_{\ell})$ of $M_{\beta,\chi}$ is smooth.
For simplicity of exposition, now we set $\chi=1$ and consider $M_{\beta}\vcentcolon=M_{\beta,1}$ (see Remark 5.4 for the case of arbitrary $\chi$). As an application of Theorem 1.1, we prove the stabilization of intersection Betti numbers of $M_{\beta}$ for Enriques and bielliptic surfaces, providing new evidence for [33, Conjecture 1.3].
Theorem 1.3 (Theorem 5.3).
Suppose that $S$ is an Enriques or bielliptic surface. Fix $k\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on $S$, there exists $d(\beta_{0},k)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$ and $k$) such that
$\dim_{\mathbb{Q}}\operatorname{IH}^{k}(M_{d\beta_{0}})=b_{k}^{\infty}\quad\text{for all integer }d\geq d(\beta_{0},k),$
where $b_{k}^{\infty}$ is defined by (1.1).
@@
Although moduli spaces of sheaves on $K3$ and abelian surfaces have been extensively studied, much less is known about the cases of Enriques and bielliptic surfaces. In the bielliptic case, recent progress was made in [27].
@@
When $S$ is a generic Enriques surface and the class of $\beta$ is not divisible by $2$ in the group $\operatorname{Num}(S)$ of numerical equivalence classes of divisors on $S$, Saccà [30] proved that $M_{\beta}$ is a smooth projective Calabi–Yau variety of dimension $\beta^{2}+1$. In this case, although the odd cohomology of $M_{\beta}$ can be nonzero as shown in [30, Theorem 5.10], Theorem 1.3 implies that the $k$-th Betti number $b_{k}(M_{\beta})$ for fixed odd $k$ vanishes for sufficiently positive $\beta$. We also prove, in this setting, a refined version of Theorem 1.3 concerning the stabilization of the invariants $n_{\beta}^{i,j}$ (defined by (2.4) with $\phi$ taken to be $h:M_{\beta}\to|\beta|$) which are conjectured to refine the Gromov–Witten/Pandharipande–Thomas invariants of the local surface $\operatorname{Tot}(K_{S})$ [10, 12, 21] and are the shifted perverse Hodge numbers considered in [28].
Theorem 1.4 (Theorem 5.7).
@@
Suppose that $S$ is a generic Enriques surface. Fix $i,j\in\mathbb{Z}_{\geq 0}$. Given any ample divisor $\beta_{0}$ on $S$ such that $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ (depending on $\beta_{0}$, $i$ and $j$) such that
$n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j}\quad\text{for all integer $d\geq d(\beta_{0},i,j)$ with }2\nmid d,$
@@
Under the assumption of [28, Conjecture A], Oberdieck proved a stronger stabilization result [28, Proposition 1.1] for the shifted perverse Hodge numbers. We note that the product formula (1.2) coincides with the one in [28, Proposition 1.1]. It remains open whether our approach, which does not rely on [28, Conjecture A], can be modified to show $n_{\beta}^{i,j}=n^{i,j}_{\infty}$ if $i<\beta^{2}/4$ and $j<\beta^{2}/4-1$. An explicit expression for $d(\beta_{0},i,j)$ in Theorem 1.4 is given in (5.27).
The rest of this paper is organized as follows. In Section 2, we collect some notation, definitions, and facts that will be useful later. Section 3 is devoted to the local versality of the universal family of curves in high codimension and thus to a proof of Theorem 1.2. In Section 4, we prove Theorem 1.1 and discuss some difficulties arising in the general case. Finally, in Section 5, we study in detail the cases of Enriques and bielliptic surfaces and prove Theorems 1.3 and 1.4.
@@
A variety is an integral and separated scheme of finite type. A curve is a projective scheme of pure dimension $1$. By a point of a scheme, we always mean a closed point. All sheaves are assumed to be coherent. For a sheaf $\mathcal{E}$} on a projective scheme $X$, we denote by $h^{i}({\mathcal{E}})$ (resp. $\operatorname{ext}^{i}({\mathcal{E}},{\mathcal{E}})$) the dimension of the $i$-th sheaf cohomology $H^{i}(X,{\mathcal{E}})$ (resp. $i$-th Ext group $\operatorname{Ext}^{i}({\mathcal{E}},{\mathcal{E}})$) as a $\mathbb{C}$-vector space.
Acknowledgments
The authors thank Weite Pi and Junliang Shen for helpful discussions. The second author is grateful to her advisor Jun Li for his constant support, to János Kollár and Claire Voisin for inspiring discussions on the deformation theory of singularities of planar curves, and to Georg Oberdieck for suggesting considering bielliptic surfaces. The first author is partially supported by the Fundamental Research Funds for the Central Universities. The second author is supported by the NSFC grants (No. 12121001 and No. 12425105) and LMNS.
@@ 2. Preliminaries
This section gathers the notation and background material that will be essential later.
2.1. Relative Hilbert schemes of points
Let $\pi:{\mathcal{C}}\to|\beta|$ be the universal family of curves in $|\beta|$. For $n\in\mathbb{Z}_{\geq 0}$, denote by
$\pi^{[n]}:{\mathcal{C}}^{[n]}\to|\beta|$
the relative Hilbert scheme of $n$ points on the fibers of $\pi$. For $C\in|\beta|$, the fiber of $\pi^{[n]}$ over $C$ is the Hilbert scheme $C^{[n]}$ parametrizing 0-dimensional, length $n$ closed subschemes of $C$. Note that $\pi^{[0]}:|\beta|\to|\beta|$ is the identity and $\pi=\pi^{[1]}:{\mathcal{C}}\to|\beta|$. For any open subscheme $U\subset|\beta|$, write ${\mathcal{C}}_{U}^{[n]}$ for $(\pi^{[n]})^{-1}(U)$. The following property of $\pi^{[n]}$ is a corollary of [18, Theorem 1.1] (see also [34, Proposition 2.7]).

### v2

@@ Abstract.
In this paper, we develop a unified approach to study the intersection Betti numbers of moduli spaces of one-dimensional semistable sheaves on smooth projective surfaces. Assuming the irreducibility of such moduli spaces, we prove that their intersection Betti numbers in a certain range of degrees coincide with the stable Betti numbers of Hilbert schemes of points. As an application, for minimal surfaces of Kodaira dimension 0, we show that these intersection Betti numbers stabilize in each fixed degree, which fits into the broader context of stable cohomology for moduli spaces of sheaves. In the case of Enriques surfaces, we also prove a refined stabilization result related to Oberdieck’s conjecture on perverse Hodge numbers.
Key words and phrases:
moduli spaces of sheaves, intersection cohomology, perverse filtrations, compactified Jacobians
2020 Mathematics Subject Classification:
14D20, 14F06 (primary), 32S60 (secondary)
@@ 1. Introduction
In recent years, moduli spaces of one-dimensional sheaves on surfaces have attracted considerable attention. On the one hand, their connections with enumerative geometry and the theory of Higgs bundles have led to a number of new conjectures and results (e.g., [3, 4, 14, 20, 21, 25, 27, 35, 36, 46]). On the other hand, these moduli spaces exhibit not only asymptotic behavior analogous to that of moduli spaces of torsion-free sheaves, but also pathologies in contrast to the torsion-free case (see [42, §5]), revealing rich geometry that remains to be understood.
Let $S$ be a connected, smooth, complex projective surface with a fixed ample divisor $H$. For a nonzero effective divisor $\beta$ on $S$ and an integer $\chi$, let $M_{\beta,\chi}$ be the coarse moduli space parametrizing S-equivalence classes of one-dimensional semistable (with respect to $H$) sheaves $\mathcal{F}$} on $S$ with fixed determinant $\det({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and Euler characteristic $\chi({\mathcal{F}})=\chi$. The moduli space $M_{\beta,\chi}$ admits a Hilbert–Chow morphism
$h:M_{\beta,\chi}\to|\beta|$
to the linear system $|\beta|$, defined by taking the Fitting support. Over the open subscheme $|\beta|^{\mathrm{int}}\subset|\beta|$ parametrizing integral curves, the fibers of $h$ are compactified Jacobians.
A natural step toward understanding the geometry of $M_{\beta,\chi}$ is to determine its topological invariants. Since $M_{\beta,\chi}$ is singular in general, we consider its intersection cohomology, which is better-behaved than the ordinary cohomology. In this paper, we study the intersection Betti numbers of $M_{\beta,\chi}$, defined as the dimensions of the intersection cohomology groups $I\!H^{k}(M_{\beta,\chi})$ (with $\mathbb{Q}$}-coefficients) for $k\in\mathbb{Z}_{\geq 0}$. Our first main result expresses these intersection Betti numbers, in a certain range of degrees, as the coefficients $b_{k}^{\infty}$ of the formal power series
(1.1) $\sum_{k\geq 0}b_{k}^{\infty}q^{k}=\prod_{m\geq 1}\frac{(1+q^{2m-1})^{b_{1}(S)}(1+q^{2m+1})^{b_{1}(S)}}{(1-q^{2m})^{b_{2}(S)+1}(1-q^{2m+2})},$
where $b_{i}(S)$ denotes the $i$-th Betti number of $S$.
@@
Suppose that $M_{\beta,\chi}$ is irreducible, and denote by
@@
the codimension of $M_{\beta,\chi}\setminus h^{-1}(|\beta|^{\mathrm{int}})$ in $M_{\beta,\chi}$. If
@@
    $\beta$ is $\max\{1,2k-2\}$-very ample (see Definition 2.4), and
@@
    $k\leq\min\left\{c-1,\frac{2}{3}\dim|\beta|\right\}$,
then
$\dim_{\mathbb{Q}}I\!H^{k}(M_{\beta,\chi})=b_{k}^{\infty}.$
By Göttsche’s formula [11], the number $b_{k}^{\infty}$ is exactly the $k$-th Betti number $b_{k}(S^{[n]})$ of the Hilbert scheme $S^{[n]}$ of $n$ points on $S$ for any integer $n\geq k$. The connection between (intersection, virtual, or ordinary) Betti numbers of moduli spaces of sheaves on $S$ and $b_{k}^{\infty}$ has been observed in various settings. For some moduli spaces of torsion-free sheaves, this connection has been established, for instance, in [8, 23]. Similar results have been obtained for moduli spaces of one-dimensional sheaves on specific surfaces such as K3 surfaces [29], the projective plane [46], and del Pezzo surfaces [36]. The above Theorem 1.1 provides a unifying perspective on the intersection Betti numbers of moduli spaces of one-dimensional sheaves on surfaces.
To prove Theorem 1.1, we adapt the approach in Sections 2 and 3 of our previous joint work with Pi and Shen [36] to a more general setting. For del Pezzo (as well as K3 and Abelian) surfaces, the open subscheme $h^{-1}(|\beta|^{\mathrm{int}})$ of $M_{\beta,\chi}$ is smooth. The main difficulty in generalizing the method in [36] to other cases arises from the singularities in $h^{-1}(|\beta|^{\mathrm{int}})$. Unlike in the torsion-free case, dimension bounds for the singular locus of $M_{\beta,\chi}$ were largely unknown, since standard techniques for torsion-free sheaves, such as those in [16, Ch. 9], are generally inapplicable to torsion sheaves. We use the deformation theory of curve singularities to show that the dimension of the singular locus of $h^{-1}(|\beta|^{\mathrm{int}})$ is governed by the positivity of $\beta$, which may be of independent interest and will be essential to our proof of Theorem 1.1. The precise statement is as follows.
@@
If $\beta$ is $\max\{1,2k-2\}$-very ample, then there exists an open subscheme $U_{k}\subset|\beta|^{\mathrm{int}}$ such that
@@
    $\operatorname{codim}(|\beta|^{\mathrm{int}}\setminus U_{k},|\beta|^{\mathrm{int}})\geq k+1$;
@@
    the open subscheme $h^{-1}(U_{k})$ of $M_{\beta,\chi}$ is smooth.
Motivated by Theorem 1.1 and the results for torsion-free sheaves, we propose the following conjecture on the stabilization of intersection Betti numbers, which can be viewed as a rank 0 analogue of [8, Conjecture 1.1].
Conjecture 1.3.
Given any ample divisor $\beta_{0}$ on $S$ and any $k\in\mathbb{Z}_{\geq 0}$, $\chi\in\mathbb{Z}$, there exists $d(\beta_{0},k,\chi)\in\mathbb{Z}_{>0}$ such that for every integer $d\geq d(\beta_{0},k,\chi)$, the intersection Betti number $\dim_{\mathbb{Q}}I\!H^{k}(M_{d\beta_{0},\chi})$ is independent of $d$. If furthermore $M_{d\beta_{0},\chi}$ is irreducible, then the stable intersection Betti number is given by
(1.2) $\dim_{\mathbb{Q}}I\!H^{k}(M_{d\beta_{0},\chi})=b_{k}^{\infty}.$
As an application of Theorem 1.1, we prove that Conjecture 1.3 holds for all minimal surfaces of Kodaira dimension 0, namely, K3, Abelian, Enriques, and bielliptic surfaces. We focus primarily on the Enriques and bielliptic cases.
Theorem 1.4.
Suppose that $S$ is a minimal surface of Kodaira dimension 0. Then Conjecture 1.3 is true. Moreover, $d(\beta_{0},k,\chi)$ can be chosen independently of $\chi$ such that (1.2) holds for all integers $d\geq d(\beta_{0},k)\vcentcolon=d(\beta_{0},k,\chi)$.
Remark 1.5.
When $S$ is a K3 or an Abelian surface, the moduli space $M_{\beta,1}$ is smooth, and the statement of Conjecture 1.3 for $\chi=1$ can be deduced from known results in the literature, such as those in [45]. In the K3 case, Mauri and Migliorini [29] proved the stabilization of intersection Betti and intersection Hodge numbers of possibly singular $M_{\beta,\chi}$. The Abelian case can be proved similarly using the argument in [29, §10.3]. As in [29, 36], an analysis of $\operatorname{codim}(|d\beta_{0}|\setminus|d\beta_{0}|^{\mathrm{int}},|d\beta_{0}|)$ is needed for our proof of Theorem 1.4. However, our approach to determining the stable intersection Betti numbers is different from that in [29] and, in particular, does not involve the deformation equivalence of moduli spaces with Hilbert schemes of points on $S$. Such equivalence fails in the Enriques case (since in this case $M_{\beta,\chi}$ is odd-dimensional, e.g., [37]) and is unknown in the bielliptic case.
@@
Although moduli spaces of sheaves on K3 and Abelian surfaces have been extensively studied, much less is known about the cases of Enriques and bielliptic surfaces. In the bielliptic case, some progress was made in [34].
@@
When $S$ is a generic Enriques surface, assuming that $\beta^{2}>0$ and the class of $\beta$ is not divisible by $2$ in the group $\operatorname{Num}(S)$ of numerical equivalence classes of divisors on $S$, Saccà [37] proved that $M_{\beta}\vcentcolon=M_{\beta,1}$ is a smooth projective Calabi–Yau variety of dimension $\beta^{2}+1$ and computed its first few Betti numbers. In this case, although the odd cohomology of $M_{\beta}$ can be nonzero as shown in [37, Theorem 5.10], our Theorem 1.4 implies that for fixed odd $k$, the $k$-th Betti number $b_{k}(M_{\beta})$ vanishes if $\beta$ is sufficiently positive. We also prove, in this setting, a refined version (Theorem 1.6 below) of Theorem 1.4 on the stabilization of the invariants $n_{\beta}^{i,j}$ (defined by (2.6) with $\varphi$ taken to be $h:M_{\beta}\to|\beta|$). These invariants are conjectured to refine the Gromov–Witten/Pandharipande–Thomas invariants of the local surface $\operatorname{Tot}(K_{S})$ [15, 18, 27] and are the shifted perverse Hodge numbers considered in [35].
Theorem 1.6.
@@
Suppose that $S$ is a generic Enriques surface. Given any ample divisor $\beta_{0}$ on $S$ with $2\nmid\beta_{0}$ in $\operatorname{Num}(S)$ and any $i,j\in\mathbb{Z}_{\geq 0}$, there exists $d(\beta_{0},i,j)\in\mathbb{Z}_{>0}$ such that for all odd integers $d\geq d(\beta_{0},i,j)$,
$n_{d\beta_{0}}^{i,j}=n_{\infty}^{i,j},$
@@
Under the assumption of [35, Conjecture A], Oberdieck proved a stronger stabilization result [35, Proposition 1.1] for the shifted perverse Hodge numbers. We note that the product formula (1.3) coincides with the one in [35, Proposition 1.1]. It remains open whether our approach, which does not rely on [35, Conjecture A], can be modified to show $n_{\beta}^{i,j}=n^{i,j}_{\infty}$ if $i<\beta^{2}/4$ and $j<\beta^{2}/4-1$. An explicit expression for $d(\beta_{0},i,j)$ in Theorem 1.6 will be given in (5.24).
The rest of this paper is organized as follows. In Section 2, we collect some notation, definitions, and results needed later. Section 3 is devoted to the local versality of the universal family of curves in high codimension and contains the proof of Theorem 1.2. In Section 4, we prove Theorem 1.1 and discuss some difficulties arising in the general case. Finally, in Section 5, we study in detail the cases of minimal surfaces of Kodaira dimension 0 to prove Theorems 1.4 and 1.6, concluding with a discussion of related open questions.
@@
All schemes are assumed to be of finite type over the field $\mathbb{C}$ of complex numbers. A variety is an integral and separated scheme. A curve is a projective scheme of pure dimension $1$. A surface is a variety of dimension $2$. By a point of a scheme, we always mean a closed point. All sheaves are assumed to be coherent. For a sheaf $\mathcal{E}$} on a projective scheme $X$, we denote by $h^{i}({\mathcal{E}})$ (resp. $\operatorname{ext}^{i}({\mathcal{E}},{\mathcal{E}})$) the dimension of the $i$-th sheaf cohomology $H^{i}(X,{\mathcal{E}})$ (resp. $i$-th Ext group $\operatorname{Ext}^{i}({\mathcal{E}},{\mathcal{E}})$) as a $\mathbb{C}$-vector space.
@@ 2. Preliminaries
2.1. Moduli spaces of one-dimensional sheaves
We first review some basic notions concerning the moduli spaces studied in this paper and then discuss some properties regarding the Hilbert–Chow morphism and irreducibility.
Fix a smooth polarized surface $(S,H)$ as in Section 1. Recall that the dimension of a sheaf is the dimension of its support. A one-dimensional sheaf $\mathcal{F}$} on $S$ is called semistable (resp. stable) (with respect to $H$) if every nonzero subsheaf $\mathcal{G}$$\subsetneq{\mathcal{F}}$ is one-dimensional with
$\frac{\chi({\mathcal{G}})}{c_{1}({\mathcal{G}})\cdot H}\leq(\text{resp.}<)\,\frac{\chi({\mathcal{F}})}{c_{1}({\mathcal{F}})\cdot H}.$
Two semistable sheaves are called S-equivalent if their Jordan–Hölder filtrations have isomorphic factors up to reordering. Each S-equivalence class is represented by a unique polystable sheaf, i.e., a semistable sheaf which is the direct sum of stable sheaves. In general, the coarse moduli space $M_{\beta,\chi}$ considered in this paper depends on the ample divisor $H$, but we have omitted $H$ from the notation since $H$ is fixed throughout.
Let ${\mathcal{M}}_{\beta,\chi}$ be the moduli stack of one-dimensional semistable sheaves on $S$ with $\det({\mathcal{F}})\cong{\mathcal{O}}_{S}(\beta)$ and $\chi({\mathcal{F}})=\chi$. Choose a large integer $n$ such that for any semistable sheaf $\mathcal{F}$} in ${\mathcal{M}}_{\beta,\chi}$, the sheaf ${\mathcal{F}}(nH)$ is globally generated and $h^{j}({\mathcal{F}}(nH))=0$ for $j>0$. Set $N\vcentcolon=(\beta\cdot H)n+\chi$. Denote by $Q$ the Quot scheme parametrizing quotients ${\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{E}}$, where $\mathcal{E}$} is semistable with fixed Hilbert polynomial $\chi({\mathcal{E}}(mH))=(\beta\cdot H)m+\chi$ and the induced map $H^{0}(S,{\mathcal{O}}_{S}^{\oplus N})\to H^{0}(S,{\mathcal{E}}(nH))$ is an isomorphism.
Let $\operatorname{Pic}(S)$ denote the Picard scheme of $S$. There is a morphism
$\det:Q\to\operatorname{Pic}(S)$
sending ${\mathcal{O}}_{S}(-nH)^{\oplus N}\to{\mathcal{E}}]$ to $\det({\mathcal{E}})$. Then ${\mathcal{M}}_{\beta,\chi}$ has a quotient stack presentation
${\mathcal{M}}_{\beta,\chi}\simeq[Q_{\beta}/\operatorname{GL}(N)],$
where $Q_{\beta}\vcentcolon=\det^{-1}({\mathcal{O}}_{S}(\beta))$ and $\operatorname{GL}(N)$ is the general linear group acting naturally on $Q_{\beta}$.
The natural morphism
$\nu:{\mathcal{M}}_{\beta,\chi}\to M_{\beta,\chi}$
is the good moduli space for ${\mathcal{M}}_{\beta,\chi}$. Denote by
$\tilde{h}=h\circ\nu:{\mathcal{M}}_{\beta,\chi}\to|\beta|$
the stacky Hilbert–Chow morphism. Since the fibers of $h$ over the locus $|\beta|^{\mathrm{int}}$ of integral curves are compactified Jacobians and have dimension $p_{a}(\beta)$, where $p_{a}(\beta)$ is the arithmetic genus of $\beta$, the fibers of $\tilde{h}$ over $|\beta|^{\mathrm{int}}$ have dimension $p_{a}(\beta)-1$. Whether the fiber dimension of $\tilde{h}$ jumps depends on numerical properties of the canonical divisor $K_{S}$.
Lemma 2.1 ([46, Corollary 1.3]).
If $K_{S}$ is nef and $|\beta|^{\mathrm{int}}\neq\varnothing$, then for any curve $C\in|\beta|$, the fiber $\tilde{h}^{-1}(C)$ has dimension $p_{a}(\beta)-1$.
This property of $\tilde{h}$ gives constraints on the geometry of $M_{\beta,\chi}$. The next proposition generalizes [25, Theorem 2.3] and [46, Theorem 1.5] beyond the del Pezzo case.